# Multi-Task Models Adversarial Attacks

### v3

@@ Appendix D More Experimental Results
We also show the evaluation results with perturbation bound in Tables 8 and 9 for NYUv2 and Tables 10 and 11 for Taskonomy.
Model Index | #Params (M) | FGSM | PGD | ||||||||||
Segm | Norm | Dept | Total | SignTotal | GB-MTA | Segm | Norm | Dept | Total | SignTotal | GB-MTA | ||
14 | 63.6 | 21.72 | 27.07 | 35.96 | 38.51 | 46.54 | 46.69 | 29.54 | 44.67 | 63.07 | 63.09 | 92.95 | 98.9 |
4 | 63.6 | 21.12 | 17.09 | 39.01 | 38.32 | 43.45 | 44.22 | 27.92 | 21.74 | 63.88 | 59.03 | 72.88 | 79.05 |
9 | 63.6 | 26.2 | 21.63 | 38.79 | 37.72 | 40.91 | 43.21 | 33.56 | 19.35 | 83.54 | 70.83 | 85.73 | 92.65 |
23 | 63.59 | 26.18 | 23.64 | 40.69 | 37.52 | 41.52 | 43.8 | 33.2 | 26.47 | 80.66 | 65.38 | 84.67 | 90.43 |
5 | 62.48 | 23.19 | 22.05 | 39.14 | 38.54 | 39.67 | 42.99 | 29.12 | 30 | 70.99 | 61.4 | 71.46 | 79.6 |
10 | 62.48 | 26.1 | 19.72 | 39.59 | 34.76 | 39.99 | 41.57 | 37.65 | 21.7 | 97.24 | 60.91 | 86.69 | 93.55 |
28 | 62.47 | 26.12 | 28.27 | 40.01 | 38.67 | 42.65 | 44.93 | 33.36 | 49.64 | 89.3 | 72.06 | 101.31 | 107.59 |
35 | 62.25 | 25.12 | 24.02 | 37.27 | 33.53 | 36.93 | 39.73 | 36.46 | 32.54 | 98.11 | 68.3 | 91.24 | 100.53 |
38 | 62.25 | 25.94 | 26.21 | 35.87 | 34.49 | 37.22 | 39.52 | 34.42 | 43.02 | 83.04 | 63.46 | 87.27 | 93.76 |
41 | 61.13 | 26.08 | 26.38 | 40.88 | 36.12 | 39.79 | 42.64 | 38.06 | 46.02 | 100.41 | 67.51 | 95.2 | 103.03 |
11 | 55.66 | 28.75 | 20.65 | 39.53 | 35.69 | 39.38 | 41.37 | 38.52 | 18.99 | 93.81 | 57.74 | 82.58 | 89.68 |
21 | 55.65 | 25.27 | 23.64 | 42.66 | 39.93 | 40.26 | 44.65 | 31.45 | 39.55 | 79.1 | 65 | 76.7 | 84.81 |
33 | 55.43 | 27.67 | 25.99 | 38.89 | 36.26 | 36.67 | 40.35 | 33.67 | 35.08 | 87.47 | 65.02 | 77.33 | 87.08 |
36 | 55.43 | 27.8 | 29.84 | 37.76 | 34.08 | 39.49 | 41.36 | 38.36 | 40.88 | 96.69 | 57.78 | 91.97 | 99.72 |
39 | 55.43 | 29.97 | 30.95 | 40.68 | 39.95 | 41.62 | 44.76 | 36.03 | 45.68 | 83.78 | 67.83 | 90.4 | 96.63 |
44 | 54.32 | 25.99 | 25.9 | 39.41 | 34.13 | 37.99 | 40.86 | 40.48 | 43.64 | 104.05 | 63.19 | 93.5 | 100.56 |
42 | 54.32 | 28.27 | 29.47 | 38.55 | 36.15 | 39.21 | 41.42 | 36.9 | 52.1 | 91.33 | 60.92 | 89.17 | 95.79 |
48 | 47.5 | 29.01 | 28.54 | 38.05 | 34.44 | 36.5 | 38.38 | 39.88 | 48.02 | 92.03 | 58.34 | 84.81 | 90.34 |
26 | 42.54 | 25.47 | 23.76 | 38.8 | 37.75 | 38.87 | 41.97 | 34.19 | 45.39 | 80.94 | 69.39 | 88.58 | 93.88 |
17 | 42.54 | 25.5 | 25.78 | 41.79 | 39.93 | 39.61 | 44.12 | 32.21 | 41.36 | 79.58 | 66.8 | 74.34 | 84.75 |
34 | 42.32 | 27.11 | 24.73 | 35.06 | 31.75 | 34.69 | 36.48 | 41.27 | 30.49 | 86.88 | 52.77 | 77 | 82.03 |
43 | 41.21 | 28.1 | 27.35 | 39.21 | 34.18 | 37.55 | 39.68 | 42.74 | 44.65 | 104.83 | 58.03 | 90.72 | 95.97 |
49 | 34.39 | 28.23 | 28.02 | 39.85 | 34.31 | 37.66 | 39.55 | 44.96 | 60.65 | 119.52 | 59.28 | 98.39 | 106.54 |
Model Index | #Params (M) | AutoAttack | Auto-SAGE | ||||||
Segm | Norm | Dept | Total | SignTotal | GB-MTA | Baseline | GB-MTA | ||
IND | 63.83 | 30.74 | 23.52 | 75.62 | 58.3 | 69.29 | 87.58 | 71.07 | 88.53 |
14 | 63.6 | 32.43 | 52.73 | 80.01 | 73.77 | 95.69 | 118.63 | 94.14 | 118.97 |
4 | 63.6 | 30.62 | 30.53 | 76.81 | 67.7 | 73.99 | 93.73 | 77.63 | 94.41 |
9 | 63.6 | 37.53 | 25.86 | 100.49 | 80.34 | 88.5 | 108.6 | 102.27 | 110.24 |
23 | 63.59 | 37 | 32.9 | 98.26 | 73.23 | 87.13 | 107.08 | 102.43 | 110.87 |
5 | 62.48 | 32.06 | 39.25 | 88.03 | 70.34 | 73.05 | 94.44 | 84.51 | 95.7 |
10 | 62.48 | 41.62 | 28.69 | 122.36 | 67.27 | 89.72 | 113.78 | 110.51 | 119.37 |
28 | 62.47 | 37.28 | 59.37 | 113.04 | 83.65 | 105.58 | 130.93 | 118.33 | 133.16 |
35 | 62.25 | 41.01 | 41.71 | 119.52 | 78.1 | 95.22 | 121.3 | 124.97 | 131.56 |
38 | 62.25 | 38.3 | 54.74 | 106.5 | 71.85 | 90.27 | 113.04 | 105.82 | 117.23 |
41 | 61.13 | 42.09 | 59.54 | 120.39 | 75.42 | 99.05 | 122.5 | 113.15 | 126.94 |
11 | 55.66 | 43.1 | 24.76 | 112.77 | 63.12 | 84.97 | 106.21 | 107.85 | 115.3 |
21 | 55.65 | 34.8 | 57.48 | 96.28 | 74.13 | 78.75 | 100.94 | 92.91 | 102.72 |
33 | 55.43 | 37.37 | 44.63 | 110.04 | 74.07 | 79.81 | 104.87 | 109.74 | 112.81 |
36 | 55.43 | 43 | 50.84 | 117.13 | 63.6 | 95.66 | 119.98 | 112.12 | 127.46 |
39 | 55.43 | 39.81 | 56.61 | 106.07 | 76.16 | 94.08 | 115.62 | 108.61 | 118.71 |
44 | 54.32 | 45.29 | 54.94 | 128.01 | 68.93 | 92.54 | 120.52 | 125.6 | 130.93 |
42 | 54.32 | 41.34 | 67.03 | 116.01 | 67.46 | 92.41 | 114.74 | 113.28 | 120.97 |
48 | 47.5 | 44.42 | 61.56 | 111.9 | 62.92 | 84.19 | 106.66 | 107.63 | 114.49 |
26 | 42.54 | 38.32 | 58.72 | 101.69 | 80.18 | 92.38 | 115.79 | 107.18 | 118.67 |
17 | 42.54 | 35.91 | 51.57 | 98.39 | 75.5 | 76.19 | 98.4 | 100.51 | 103.25 |
34 | 42.32 | 45.56 | 38.38 | 108.03 | 55.82 | 79.52 | 96.55 | 95.69 | 105.38 |
43 | 41.21 | 47.44 | 59.7 | 131.97 | 61.11 | 89.67 | 115.1 | 122.87 | 128.99 |
49 | 34.39 | 48.4 | 70.91 | 136.53 | 61.73 | 92.81 | 119.81 | 125.91 | 130.8 |
AS | 21.28 | 56.39 | 69.6 | 133.54 | 67.42 | 88.1 | 108.57 | 112.56 | 119.22 |
Model Index | #Params (M) | FGSM | PGD | ||||||||||||
Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | ||
190 | 105.03 | 73.19 | 23.13 | 42.67 | 13.75 | 16.55 | 72.87 | 74.15 | 190.91 | 49.41 | 138.67 | 14.98 | 14.02 | 232.55 | 236.63 |
348 | 105.02 | 67.60 | 23.05 | 49.61 | 14.74 | 14.86 | 70.57 | 78.14 | 202.55 | 51.07 | 147.26 | 19.62 | 14.74 | 246.88 | 249.54 |
200 | 104.81 | 67.83 | 19.17 | 44.49 | 12.39 | 14.55 | 71.01 | 76.78 | 182.76 | 50.02 | 139.60 | 13.13 | 16.56 | 226.76 | 233.80 |
352 | 104.80 | 67.17 | 22.59 | 44.80 | 15.19 | 14.85 | 68.88 | 74.70 | 178.64 | 48.81 | 143.03 | 17.71 | 13.70 | 224.02 | 234.34 |
469 | 104.57 | 72.08 | 22.52 | 49.17 | 17.73 | 15.17 | 73.25 | 79.48 | 182.49 | 49.82 | 155.29 | 27.55 | 18.20 | 228.44 | 240.80 |
358 | 103.68 | 69.37 | 23.91 | 46.00 | 15.62 | 15.60 | 69.91 | 74.85 | 189.90 | 50.42 | 152.70 | 18.14 | 15.53 | 237.12 | 247.29 |
481 | 103.46 | 72.21 | 22.81 | 46.38 | 21.68 | 16.24 | 74.59 | 81.94 | 212.18 | 51.14 | 157.02 | 38.94 | 24.78 | 256.71 | 266.46 |
191 | 98.21 | 68.01 | 20.88 | 40.81 | 12.19 | 13.79 | 68.19 | 69.76 | 190.28 | 52.63 | 153.36 | 13.94 | 12.94 | 226.54 | 231.65 |
959 | 96.86 | 69.48 | 23.28 | 44.76 | 21.98 | 18.80 | 73.29 | 76.47 | 201.59 | 48.17 | 163.56 | 15.56 | 17.64 | 241.75 | 256.35 |
958 | 83.75 | 74.68 | 23.46 | 47.23 | 22.66 | 16.49 | 80.55 | 84.18 | 210.33 | 50.29 | 166.25 | 15.22 | 13.20 | 250.88 | 266.38 |
1020 | 83.53 | 74.85 | 22.43 | 48.32 | 20.34 | 16.31 | 80.78 | 86.30 | 222.92 | 49.98 | 139.97 | 18.25 | 18.01 | 259.27 | 263.03 |
1043 | 75.59 | 77.23 | 21.55 | 53.74 | 22.06 | 23.22 | 84.89 | 94.04 | 205.27 | 44.26 | 145.47 | 23.99 | 38.96 | 246.90 | 250.79 |
1037 | 62.48 | 72.67 | 22.63 | 48.99 | 22.62 | 17.82 | 77.62 | 84.68 | 216.15 | 45.51 | 152.47 | 17.61 | 47.35 | 252.55 | 261.82 |
Model Index | #Params (M) | AutoAttack | Auto-SAGE | |||||||
Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | Baseline | GB-MTA | ||
190 | 105.03 | 175.02 | 45.06 | 130.69 | 14.26 | 13.64 | 213.47 | 218.29 | 216.32 | 218.88 |
348 | 105.02 | 184.25 | 46.19 | 137.71 | 18.80 | 13.90 | 225.45 | 229.03 | 228.10 | 229.64 |
200 | 104.81 | 168.93 | 44.77 | 130.06 | 12.55 | 15.74 | 208.48 | 215.48 | 210.54 | 216.37 |
352 | 104.80 | 164.72 | 44.09 | 134.15 | 16.84 | 13.15 | 205.59 | 216.01 | 207.59 | 216.50 |
469 | 104.57 | 168.58 | 44.55 | 143.19 | 26.30 | 17.31 | 210.28 | 220.91 | 212.21 | 221.93 |
358 | 103.68 | 174.80 | 46.32 | 142.73 | 17.07 | 14.73 | 217.99 | 227.61 | 220.93 | 228.97 |
481 | 103.46 | 193.75 | 45.87 | 146.31 | 37.71 | 21.32 | 234.60 | 244.78 | 236.99 | 246.22 |
191 | 98.21 | 174.37 | 47.96 | 142.62 | 13.05 | 12.25 | 208.24 | 213.53 | 211.71 | 214.30 |
959 | 96.86 | 183.75 | 44.46 | 150.37 | 14.70 | 17.04 | 221.31 | 234.58 | 223.54 | 234.56 |
958 | 83.75 | 192.26 | 45.68 | 152.92 | 14.37 | 12.80 | 229.57 | 242.97 | 232.22 | 243.31 |
1020 | 83.53 | 202.50 | 44.34 | 130.95 | 17.02 | 16.89 | 235.80 | 240.26 | 238.95 | 241.14 |
1043 | 75.59 | 189.32 | 39.97 | 136.29 | 22.55 | 36.50 | 227.23 | 233.25 | 230.54 | 234.28 |
1037 | 62.48 | 197.99 | 41.05 | 141.10 | 16.88 | 40.50 | 230.86 | 240.70 | 240.00 | 246.43 |
Model Index | #Params (M) | FGSM | PGD | ||||||||||
Segm | Norm | Dept | Total | SignTotal | GB-MTA | Segm | Norm | Dept | Total | SignTotal | GB-MTA | ||
14 | 63.60 | 19.73 | 10.82 | 33.63 | 33.88 | 36.30 | 39.51 | 27.67 | 12.33 | 64.64 | 57.08 | 63.67 | 73.42 |
4 | 63.60 | 17.57 | 20.65 | 27.17 | 33.05 | 40.21 | 41.11 | 25.28 | 35.17 | 42.23 | 47.76 | 66.83 | 73.81 |
9 | 63.60 | 16.99 | 9.37 | 29.26 | 32.13 | 35.28 | 37.64 | 23.91 | 13.30 | 44.03 | 44.93 | 51.26 | 58.82 |
23 | 63.59 | 19.59 | 13.92 | 32.88 | 32.01 | 35.67 | 38.01 | 27.56 | 17.54 | 61.19 | 52.80 | 62.37 | 70.10 |
5 | 62.48 | 17.74 | 12.82 | 30.67 | 32.45 | 33.11 | 37.15 | 24.51 | 17.33 | 48.98 | 46.82 | 51.18 | 59.36 |
10 | 62.48 | 21.43 | 10.35 | 34.04 | 30.76 | 35.13 | 37.51 | 31.32 | 13.71 | 73.02 | 50.90 | 64.28 | 73.08 |
28 | 62.47 | 20.33 | 20.16 | 33.82 | 35.17 | 39.46 | 41.95 | 27.68 | 35.74 | 63.97 | 56.20 | 73.24 | 80.54 |
35 | 62.25 | 20.25 | 18.54 | 30.89 | 30.84 | 34.10 | 37.00 | 27.91 | 29.93 | 58.75 | 50.02 | 64.14 | 71.00 |
38 | 62.25 | 19.56 | 15.72 | 33.75 | 31.01 | 34.50 | 37.92 | 29.23 | 22.82 | 74.39 | 54.53 | 66.51 | 76.85 |
41 | 61.13 | 21.00 | 19.33 | 34.84 | 32.15 | 36.69 | 39.67 | 31.10 | 32.23 | 75.81 | 54.53 | 71.75 | 79.93 |
11 | 55.66 | 22.12 | 10.87 | 33.89 | 30.51 | 33.73 | 36.88 | 31.73 | 12.76 | 70.28 | 47.48 | 60.24 | 68.96 |
21 | 55.65 | 18.96 | 15.58 | 34.51 | 34.17 | 34.54 | 39.08 | 26.53 | 23.81 | 56.20 | 50.44 | 56.26 | 64.85 |
33 | 55.43 | 21.55 | 18.91 | 31.83 | 29.54 | 34.62 | 37.13 | 31.01 | 28.48 | 72.41 | 47.68 | 66.77 | 74.71 |
36 | 55.43 | 21.64 | 21.18 | 33.84 | 34.08 | 36.15 | 39.67 | 29.49 | 32.03 | 62.09 | 54.67 | 67.51 | 74.64 |
39 | 55.43 | 19.99 | 16.21 | 31.60 | 30.32 | 31.06 | 35.11 | 27.57 | 24.14 | 62.35 | 50.84 | 57.22 | 66.48 |
44 | 54.32 | 20.61 | 19.00 | 34.79 | 30.57 | 35.03 | 38.05 | 33.02 | 30.36 | 76.81 | 52.04 | 69.01 | 76.70 |
42 | 54.32 | 20.61 | 20.92 | 32.00 | 30.52 | 34.49 | 37.13 | 30.17 | 36.93 | 66.13 | 49.23 | 66.79 | 73.41 |
48 | 47.50 | 22.34 | 20.88 | 33.26 | 29.65 | 33.05 | 35.71 | 32.75 | 33.81 | 70.45 | 48.91 | 65.49 | 71.34 |
26 | 42.54 | 19.16 | 17.23 | 32.22 | 33.51 | 35.20 | 38.65 | 28.21 | 32.11 | 58.06 | 54.17 | 63.98 | 71.52 |
17 | 42.54 | 19.70 | 17.81 | 34.21 | 35.10 | 34.35 | 39.69 | 26.89 | 28.06 | 57.05 | 52.25 | 56.21 | 65.35 |
34 | 42.32 | 22.45 | 15.94 | 30.89 | 28.23 | 31.25 | 33.59 | 33.91 | 21.11 | 64.29 | 44.38 | 57.97 | 63.58 |
43 | 41.21 | 22.92 | 19.65 | 34.96 | 30.41 | 34.54 | 37.20 | 35.07 | 30.38 | 78.53 | 49.62 | 68.22 | 74.39 |
49 | 34.39 | 22.75 | 21.15 | 35.46 | 30.18 | 34.34 | 36.90 | 37.63 | 43.02 | 86.58 | 51.52 | 73.98 | 81.24 |
Model Index | #Params (M) | APGD | Auto-SAGE | ||||||
Segm | Norm | Dept | Total | SignTotal | GB-MTA | Baseline | GB-MTA | ||
IND | 63.83 | 27.16 | 15.28 | 48.07 | 43.27 | 46.68 | 62.72 | 50.60 | 63.78 |
14 | 63.60 | 31.53 | 15.75 | 77.52 | 65.34 | 64.94 | 86.42 | 80.49 | 88.43 |
4 | 63.60 | 28.65 | 41.38 | 51.30 | 54.93 | 67.88 | 88.39 | 66.46 | 87.21 |
9 | 63.60 | 27.10 | 17.16 | 52.03 | 51.10 | 51.84 | 69.28 | 58.22 | 70.20 |
23 | 63.59 | 31.31 | 21.70 | 74.76 | 59.86 | 63.49 | 83.53 | 78.71 | 85.70 |
5 | 62.48 | 27.91 | 21.99 | 59.31 | 54.01 | 51.87 | 70.46 | 62.59 | 71.73 |
10 | 62.48 | 35.73 | 17.45 | 91.60 | 57.20 | 65.46 | 88.04 | 81.73 | 92.48 |
28 | 62.47 | 31.55 | 43.78 | 80.85 | 65.45 | 74.95 | 98.42 | 84.99 | 97.51 |
35 | 62.25 | 32.18 | 37.56 | 72.78 | 57.23 | 65.31 | 85.37 | 74.98 | 85.89 |
38 | 62.25 | 34.16 | 28.68 | 92.92 | 63.05 | 68.05 | 94.19 | 92.69 | 99.99 |
41 | 61.13 | 35.40 | 40.66 | 93.92 | 62.34 | 73.43 | 96.80 | 86.64 | 100.73 |
11 | 55.66 | 36.47 | 16.01 | 87.30 | 53.81 | 61.46 | 83.26 | 80.89 | 88.73 |
21 | 55.65 | 29.81 | 32.22 | 68.68 | 57.18 | 57.15 | 76.41 | 69.18 | 78.34 |
33 | 55.43 | 35.83 | 35.63 | 90.57 | 54.05 | 68.32 | 92.08 | 83.96 | 96.00 |
36 | 55.43 | 33.51 | 39.54 | 77.22 | 61.78 | 68.97 | 88.82 | 80.06 | 89.48 |
39 | 55.43 | 31.24 | 29.83 | 78.41 | 58.07 | 58.41 | 79.38 | 78.94 | 83.61 |
44 | 54.32 | 38.22 | 37.76 | 97.08 | 58.90 | 67.84 | 94.07 | 90.86 | 98.39 |
42 | 54.32 | 34.23 | 45.68 | 84.02 | 55.78 | 68.21 | 88.75 | 80.95 | 89.80 |
48 | 47.50 | 37.40 | 43.53 | 86.64 | 54.54 | 64.54 | 85.45 | 81.61 | 89.60 |
26 | 42.54 | 32.45 | 41.99 | 71.78 | 62.53 | 65.45 | 86.41 | 77.90 | 87.52 |
17 | 42.54 | 30.39 | 34.78 | 68.70 | 58.85 | 57.18 | 76.24 | 74.42 | 79.05 |
34 | 42.32 | 38.77 | 25.92 | 79.02 | 49.26 | 59.24 | 75.29 | 70.50 | 79.46 |
43 | 41.21 | 40.38 | 39.16 | 99.14 | 54.95 | 67.08 | 90.60 | 88.50 | 97.06 |
49 | 34.39 | 41.65 | 51.01 | 103.47 | 55.19 | 70.24 | 94.25 | 91.96 | 99.57 |
AS | 21.28 | 44.80 | 49.65 | 99.32 | 56.27 | 66.05 | 84.44 | 79.27 | 91.01 |
Model Index | #Params (M) | FGSM | PGD | ||||||||||||
Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | ||
190 | 105.03 | 53.91 | 11.82 | 34.50 | 5.91 | 6.65 | 56.94 | 58.89 | 105.63 | 37.71 | 116.13 | 6.74 | 6.36 | 141.97 | 157.73 |
348 | 105.02 | 53.04 | 12.33 | 40.33 | 6.35 | 5.90 | 59.07 | 64.92 | 115.81 | 38.49 | 120.48 | 9.71 | 6.22 | 153.14 | 166.49 |
200 | 104.81 | 53.40 | 11.13 | 36.40 | 5.03 | 6.38 | 56.96 | 61.64 | 104.28 | 36.37 | 113.03 | 5.74 | 7.49 | 140.94 | 155.51 |
352 | 104.80 | 51.95 | 11.90 | 35.91 | 6.08 | 5.80 | 55.51 | 59.68 | 102.05 | 36.28 | 116.72 | 8.21 | 5.88 | 138.80 | 156.63 |
469 | 104.57 | 53.07 | 12.95 | 40.12 | 9.45 | 7.10 | 58.58 | 64.07 | 105.87 | 37.13 | 122.31 | 16.36 | 9.68 | 144.30 | 160.39 |
358 | 103.68 | 53.60 | 12.46 | 36.65 | 6.14 | 5.72 | 56.95 | 60.86 | 107.96 | 38.20 | 123.47 | 8.41 | 6.16 | 148.42 | 164.93 |
481 | 103.46 | 57.14 | 13.82 | 37.24 | 11.58 | 7.93 | 60.54 | 65.79 | 121.96 | 36.53 | 125.33 | 21.67 | 14.79 | 158.55 | 175.44 |
191 | 98.21 | 51.06 | 11.31 | 33.85 | 4.79 | 5.30 | 54.51 | 56.15 | 101.20 | 41.67 | 121.75 | 5.97 | 5.29 | 136.27 | 151.80 |
959 | 96.86 | 51.77 | 12.40 | 37.55 | 8.17 | 8.17 | 59.28 | 63.09 | 113.43 | 37.06 | 127.30 | 5.54 | 7.52 | 149.44 | 170.95 |
958 | 83.75 | 56.76 | 11.56 | 38.07 | 8.34 | 6.16 | 63.94 | 66.90 | 116.17 | 37.81 | 132.04 | 6.21 | 5.85 | 153.90 | 177.06 |
1020 | 83.53 | 57.77 | 11.82 | 37.90 | 8.13 | 6.36 | 64.77 | 68.43 | 126.77 | 36.15 | 115.46 | 6.72 | 7.59 | 158.30 | 175.65 |
1043 | 75.59 | 59.98 | 11.12 | 40.90 | 8.55 | 12.49 | 67.54 | 73.56 | 119.75 | 32.93 | 116.46 | 8.93 | 29.73 | 155.25 | 170.59 |
1037 | 62.48 | 55.33 | 12.16 | 36.44 | 8.69 | 8.28 | 61.76 | 66.87 | 123.41 | 32.09 | 118.24 | 7.56 | 33.09 | 155.66 | 172.77 |
Model Index | #Params (M) | APGD | Auto-SAGE | |||||||
Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | Baseline | GB-MTA | ||
190 | 105.03 | 98.82 | 33.70 | 109.26 | 6.39 | 6.12 | 132.72 | 146.86 | 134.60 | 147.75 |
348 | 105.02 | 107.45 | 34.00 | 112.95 | 9.31 | 5.93 | 142.54 | 154.78 | 144.55 | 155.85 |
200 | 104.81 | 97.78 | 32.08 | 105.60 | 5.39 | 7.20 | 131.47 | 145.08 | 133.21 | 145.95 |
352 | 104.80 | 95.66 | 32.25 | 109.25 | 7.75 | 5.62 | 129.44 | 145.90 | 131.41 | 147.07 |
469 | 104.57 | 98.82 | 33.00 | 113.63 | 15.70 | 9.14 | 134.73 | 149.15 | 136.58 | 150.36 |
358 | 103.68 | 100.97 | 34.09 | 115.14 | 7.88 | 5.74 | 138.28 | 153.40 | 140.26 | 154.84 |
481 | 103.46 | 113.34 | 32.11 | 116.28 | 20.99 | 12.67 | 147.42 | 162.81 | 149.73 | 164.18 |
191 | 98.21 | 94.75 | 37.13 | 113.48 | 5.59 | 5.13 | 127.89 | 141.74 | 129.68 | 142.58 |
959 | 96.86 | 105.62 | 33.55 | 118.21 | 5.25 | 7.28 | 139.53 | 158.45 | 141.08 | 159.39 |
958 | 83.75 | 108.44 | 33.77 | 122.67 | 5.97 | 5.76 | 143.50 | 163.92 | 145.27 | 164.72 |
1020 | 83.53 | 117.91 | 31.86 | 107.94 | 6.30 | 7.10 | 147.29 | 162.43 | 149.50 | 163.90 |
1043 | 75.59 | 112.46 | 29.12 | 109.43 | 8.45 | 27.97 | 145.09 | 159.55 | 147.15 | 160.82 |
1037 | 62.48 | 115.08 | 28.25 | 110.43 | 7.28 | 29.00 | 145.04 | 160.81 | 151.71 | 165.31 |
Results of MTL Adversarial Transferability

### v4

@@ E.2 Attack on Clean Multi-Task Models
We also show the evaluation results with perturbation bound in Tables 12 and 13 for NYUv2 and Tables 14 and 15 for Taskonomy.
We further compare DGBA with a variant of Total called WeightedTOTAL, which uses grid search to determine optimal task-specific weights when generating adversarial samples. Specifically, each task-specific gradient is assigned a weight selected from the range 0.1 to 0.9 in increments of 0.1, with the sum of the weights constrained to 1. WeightedTOTAL is integrated with PGD and APGD for comparison against DGBA.
Table 16 reports the optimal weights and the ARP achieved by the WeightedTOTAL attack on all 25 multi-task models trained on NYUv2 with Deeplab-ResNet34, using a perturbation bound of . Overall, DGBA achieves a higher ARP in 49 out of 50 cases compared to WeightedTOTAL. Additionally, it is noteworthy that WeightedTOTAL incurs significant computational overhead due to its grid search process, while DGBA introduces negligible overhead compared to the original single-task attacks, as it retains the core attack mechanisms with only additional dynamic gradient weighting.
E.3 Black-Box Attack on Clean Multi-Task Models
We further evaluate the effectiveness of DGBA under the black-box scenarios to verify whether adversarial examples generated from a local surrogate model can be effectively transferred to a target black-box model. Specifically, the experiment is conducted on multi-task models trained on the NYUv2 dataset under the classic transfer-based black-box attack setting:
Surrogate model. We generate adversarial examples using white-box attacks on the all-shared model (i.e., AS) with a fixed perturbation budget of .
Target models. These adversarial samples are then evaluated, without any access to gradients, on 7 additional multi-task networks that vary in their parameter-sharing strategies, ranging from fully independent task models (i.e., IND) to partially shared configurations.
The results are summarized in Table 6. We highlight three key observations. First, adversarial examples generated via DGBA on the surrogate model demonstrate transferability across models with different sharing structures. In particular, DGBA achieves the highest ARP in 18 out of the 21 evaluated settings. Second, the degree of transferability is closely tied to the architectural similarity between the surrogate and target models. The lowest ARP occurs when transferring to the fully independent models (i.e., IND), which has no shared parameters and is structurally furthest from the all-shared surrogate. Finally, as expected, we observe a natural performance gap between black-box and white-box scenarios. For instance, the last row of Table 6, corresponding to the white-box attack on the surrogate model itself, shows significantly higher ARP values than the black-box transfer cases. Similar comparisons can be made between Table 6 and the Table 1 in the main paper.
| Model | #Params (M) | PGD | APGD | Auto-SAGA | ||||||||||||
| Segm | Norm | Dept | Total | SignTotal | DGBA | Segm | Norm | Dept | Total | SignTotal | DGBA | Baseline | DGBA | |||
| BB | IND | 63.83 | 23.52 | 11.21 | 20.83 | 25.57 | 22.85 | 25.14 | 23.55 | 11.38 | 21.23 | 25.65 | 22.84 | 25.32 | 25.56 | 25.14 |
| 5 | 62.48 | 24.80 | 13.07 | 24.41 | 27.06 | 24.78 | 27.21 | 24.82 | 13.16 | 24.90 | 27.25 | 24.81 | 27.46 | 27.10 | 27.19 | |
| 35 | 62.25 | 28.71 | 16.92 | 44.55 | 34.36 | 37.06 | 41.52 | 28.73 | 16.94 | 44.96 | 34.48 | 37.08 | 41.87 | 34.34 | 41.56 | |
| 41 | 61.13 | 30.62 | 18.46 | 44.77 | 36.32 | 39.01 | 43.44 | 30.77 | 18.57 | 45.35 | 36.51 | 38.92 | 43.78 | 36.24 | 43.44 | |
| 21 | 55.65 | 26.10 | 13.29 | 25.50 | 28.52 | 26.40 | 28.91 | 26.06 | 13.38 | 25.97 | 28.65 | 26.43 | 29.27 | 28.56 | 28.92 | |
| 39 | 55.43 | 26.47 | 15.63 | 28.32 | 29.38 | 28.16 | 31.45 | 26.54 | 15.73 | 28.90 | 29.63 | 28.24 | 31.83 | 29.45 | 31.43 | |
| 26 | 42.54 | 28.69 | 19.80 | 30.70 | 32.07 | 32.80 | 36.43 | 28.70 | 20.03 | 31.19 | 32.15 | 32.74 | 36.64 | 32.06 | 36.46 | |
| WB | AS | 21.28 | 37.80 | 37.87 | 78.58 | 49.49 | 64.36 | 69.60 | 44.80 | 49.65 | 99.32 | 56.27 | 66.05 | 84.44 | 79.27 | 91.01 |
E.4 Attack on Adversarially Trained Models
As introduced in Section 4.3 in the main paper, we adopt the single task Friendly Adversarial Training (FAT) [zhang2020attacks] in the MTL context. Specifically, we modify the underlying adversarial samples generation process to the multi-task attacks we investigated in this paper, including naive multi-task adaptations Single and Total, and the proposed DGBA. The detailed adversarial training algorithm is described in Algorithm E.4, where for each mini-batch, we generate the adversarial samples with DGBA-PGD to train the model.
Algorithm 2 FAT DGBA
Table 17 reports the accuracy of multi-task models trained on NYUv2 with adversarial training. It includes metrics for all tasks and performance degradation (shown in columns containing ARP). We employ PGD-based adapted multi-task attacks Single-x and Total as well as DGBA with to generate different adversarial data. It can be seen that after adversarial training, the average accuracy of the multi-task model is dropped by compared with training with clean data only (the “w/o AT” row).
The same phenomenon of decreased model accuracy and increased model robustness can also be observed in Tables 18 and 19, where FGSM-based multi-task attack variants are utilized when generating adversarial samples in adversarial training. Table 18 reports the accuracy of the adversarially trained multi-task models similar to Table 17, while Table 19 shows the ARP of multi-task models trained with and without FAT in Table 18 and attacked by multi-task attacks including FGSM variants of DGBA, Single, and Total with .
We first observe a accuracy drop with adversarial training from Table 18, that is, decreased model performance. Then we observe an increase in model robustness after adversarial training from the lower ARP after the attack reported in Table 19. For instance, when attacked by DGBA, ARP decreased from to . Furthermore, DGBA remains the most effective attack method, consistently outperforming the Single-x and Total baselines by up to 6.87%.
E.5 Parameter Sharing and Model Robustness